OpenStudy (anonymous):

Help please and thank you :)!! x1/3 times x 1/3

OpenStudy (anonymous):

\[x \frac{ 1 }{ 3 } \times x \frac{ 1 }{ 3 }\]

OpenStudy (anonymous):

I'm confused would it stay as 1/3 or 1/9?

OpenStudy (aripotta):

1/9 i think

OpenStudy (aripotta):

cuz when you multiply fractions, you multiply the numerators then the denominators. giving you 1/9

OpenStudy (aripotta):

and x times x is x^2

OpenStudy (anonymous):

Thank you lovely!!)

OpenStudy (aripotta):

no problem

OpenStudy (anonymous):

:)